# Exploring Context Window of Large Language Models via Decomposed Positional Vectors

### v1

@@ 3.1 Experimental Settings
To better analyze positional information, we consider model variants with different positional encodings (PE) and attention mechanisms: variants without positional encodings (NoPE) [19] as well as variants with two different positional encodings: RoPE [7] and ALiBi [6]. We continually pre-train the TinyLlama-1.1B checkpoint [23] on 50B tokens from RedPajama [24] with a context window $C=2048$, resulting in a set of comparison models with different positional encodings and attention mechanisms, as shown in the Table 1. Full attention means that each token can attend to all previous tokens, while window attention restricts each token to attend only to previous tokens within a window size $W$. The training details are described in Appendix B.
@@ 3.2.1 Formation of Positional Vectors with Full Attention
To study how positional information is distributed over different positions, we first visualize the positional vectors $\mathbf{p}_{1,t}$ (Eq. 4) decomposed from the outputs of the first layer using principal component analysis (PCA). As shown in Figure 2 (left column), initial tokens (e.g., $\leq 4$ tokens) exhibit significantly distinct positional vectors, while the positional vectors of subsequent tokens are similar to each other. As a comparison, we also present the PCA results of all positional vectors at the 7-th layer. Interestingly, position vectors are evenly distributed across all the positions in Figure 2 (right column). Such a finding indicates that position vectors have captured the corresponding positional information since these vectors are distinct from each other across positions. In other words, being distinct can be considered as a kind of positional evidence. By comparing the left and right columns of Figure 2, it seems that only initial tokens are different from the rest tokens after the first layer, which might suggest that after the first layer, initial tokens have already formed distinct positional information but subsequent tokens have not yet established such information. To investigate the reasons behind this phenomenon, we select the first attention head in the first layer (similar to other heads) to analyze attention scores, as detailed in Appendix C. We can prove that the positional vector $\mathbf{p}_{1,1}$ for the first token is different from the following tokens and the attention scores affect the formation of positional information. Thus, through several layers, the tokens after the first token will gradually form distinct positional vectors (Figure 2 right column).
@@ 3.3.1 Direct Extrapolation
Table 3: The interpolation results of positional vectors, where Factor ($={\text{Target Length}}/{C}$) is the expansion factor of the context window, Ratio is the effective interpolation ratio of positional vectors (detailed in Appendix D), and Similarity is the average cosine similarity between the scaled positional vector and the original most similar positional vector by averaging all layers.
@@ 4 Extending Context Window via Positional Vectors
Inspired by our analysis of the formation of positional vectors and the interpolation of positional vectors when extending the context window, we propose two training-free context window extension methods, i.e., positional vector replacement and attention window extension. The pseudocode of these methods can presented in Appendix F.
@@ 4.1 Positional Vector Replacement
Specifically, we linearly interpolate the positional vectors within the context window with an interpolation ratio $r$ and multiply the interpolated ones with a times $\alpha~{}(\geq 1)$. In practice, we find that properly increasing the interpolation ratio $r$ and times $\alpha$ can achieve better effectiveness of interpolation (details are discussed in Appendix E). Owing to the critical role of initial tokes, the positional vectors of the first four tokens remain unchanged, while those of subsequent tokens are replaced with the interpolated vectors. The replaced output $\hat{\mathbf{h}}_{l,t}$ for each layer can be formulated as:
@@ 4.3 Results on Language Modeling
To assess the effectiveness of our proposed methods, we evaluate language modeling performance on the test set of PG-19 [22]. In line with previous work [6], we measure PPL across various input lengths (from 2K to 8K) using a sliding window approach. We apply positional vector replacement to TL-NoPE and attention window extension to TL-Window. For compared baselines, we select Dynamic-NTK [11] for TL-RoPE and Attention Scaling [20] for TL-NoPE.
@@ 5 Related Work
LLMs were often constrained by pre-defined context windows. When processing inputs that exceed these windows, models typically encountered OOD issues, leading to significant performance degradation. Various methods were proposed to address this limitation and model longer texts, which can be roughly categorized into length extrapolation and context window extension [30]. Length extrapolation techniques aimed to maintain stable PPL regardless of text length by designing specialized positional encodings or window attention mechanisms [6, 17, 18, 15, 14]. Conversely, context window extension methods focused on extending the context window of existing models by adapting positional encodings or temperature hyper-parameters, thereby enlarging the context window with minimal performance loss [13, 12, 16, 20, 11, 10]. This paper bridges the concepts of length extrapolation and context window extension through the lens of positional vectors, enhancing the interpretability of context windows in LLMs.
@@ 6 Conclusion
Appendix A Limitation
Our work provides an extensive discussion and analysis of the context window through the lens of positional vectors. However, our study is constrained by the use of small-scale LLMs that we trained ourselves, due to the unavailability of existing LLMs with the specific positional encodings and attention patterns required for our experiments. Furthermore, we have demonstrated the effectiveness of our proposed methods solely on our own models, again limited by the absence of suitable external models. In future work, we aim to seek a broader range of models to validate our findings more comprehensively.
Appendix B Training and Experimental Details
@@
Appendix C The Positional Vectors After the First Layer
@@ C.1 Details
C.2 Proof of Preference in Attention Scores
@@
Appendix D Defination of Effective Interpolation Ratio
@@
Appendix E Analysis of Positional Vector Replacement
E.1 Optimal Replacement Layer
@@
E.2 Effective Interpolation Ratio
@@
Appendix F Pseudo Code
F.1 Positional Vector Replacement
@@
F.2 Attention Window Extension

### v2

@@ 3.1 Experimental Settings
To better analyze positional information, we consider model variants with different positional encodings (PE) and attention mechanisms: variants without positional encodings (NoPE) [19] as well as variants with two different positional encodings: RoPE [7] and ALiBi [6]. We continually pre-train the TinyLlama-1.1B checkpoint [23] on 50B tokens from RedPajama [24] with a context window $C=2048$, resulting in a set of comparison models with different positional encodings and attention mechanisms, as shown in the Table 1. Full attention means that each token can attend to all previous tokens, while window attention restricts each token to attend only to previous tokens within a window size $W$. The training details are described in Appendix A. We also evaluate common LLMs (e.g., Llama-3-8B) and LLMs without positional encodings trained from scratch, and the evaluated results are listed in Appendix F.
@@ 3.2.1 Formation of Positional Vectors with Full Attention
To study how positional information is distributed over different positions, we first visualize the positional vectors $\mathbf{p}_{1,t}$ (Eq. 4) decomposed from the outputs of the first layer using principal component analysis (PCA). As shown in Figure 2 (left column), initial tokens (e.g., $\leq 4$ tokens) exhibit significantly distinct positional vectors, while the positional vectors of subsequent tokens are similar to each other. As a comparison, we also present the PCA results of all positional vectors at the 7-th layer. Interestingly, position vectors are evenly distributed across all the positions in Figure 2 (right column). Such a finding indicates that position vectors have captured the corresponding positional information since these vectors are distinct from each other across positions. In other words, being distinct can be considered as a kind of positional evidence. By comparing the left and right columns of Figure 2, it seems that only initial tokens are different from the rest tokens after the first layer, which might suggest that after the first layer, initial tokens have already formed distinct positional information but subsequent tokens have not yet established such information. To investigate the reasons behind this phenomenon, we select the first attention head in the first layer (similar to other heads) to analyze attention scores, as detailed in Appendix B. We can prove that the positional vector $\mathbf{p}_{1,1}$ for the first token is different from the following tokens and the attention scores affect the formation of positional information. Thus, through several layers, the tokens after the first token will gradually form distinct positional vectors (Figure 2 right column).
@@ 3.3.1 Direct Extrapolation
Table 3: The interpolation results of positional vectors, where Factor ($={\text{Target Length}}/{C}$) is the expansion factor of the context window, Ratio is the effective interpolation ratio of positional vectors (detailed in Appendix C), and Similarity is the average cosine similarity between the scaled positional vector and the original most similar positional vector by averaging all layers.
@@ 4 Extending Context Window via Positional Vectors
Inspired by our analysis of the formation of positional vectors and the interpolation of positional vectors when extending the context window, we propose two training-free context window extension methods, i.e., positional vector replacement and attention window extension. The pseudocode of these methods can presented in Appendix E.
@@ 4.1 Positional Vector Replacement
Specifically, we linearly interpolate the positional vectors within the context window with an interpolation ratio $r$ and multiply the interpolated ones with a times $\alpha~{}(\geq 1)$. In practice, we find that properly increasing the interpolation ratio $r$ and times $\alpha$ can achieve better effectiveness of interpolation (details are discussed in Appendix D). Owing to the critical role of initial tokes, the positional vectors of the first four tokens remain unchanged, while those of subsequent tokens are replaced with the interpolated vectors. The replaced output $\hat{\mathbf{h}}_{l,t}$ for each layer can be formulated as:
@@ 4.3 Results on Language Modeling
To assess the effectiveness of our proposed methods, we evaluate language modeling performance on the test set of PG-19 [22]. In line with previous work [6], we measure PPL across various input lengths (from 2K to 8K) using a sliding window approach. We apply positional vector replacement to TL-NoPE and attention window extension to TL-Window. All the hyper-parameters are selected according to the PPL and the change of positional vectors across layers. For compared baselines, we select Dynamic-NTK [11] for TL-RoPE and Attention Scaling [20] for TL-NoPE.
@@ 5 Related Work
LLMs were often constrained by pre-defined context windows. When processing inputs that exceed these windows, models typically encountered OOD issues, leading to significant performance degradation. To meet the growing demands of long context tasks [30, 31], various methods were proposed to address this limitation and model longer texts, which can be roughly categorized into length extrapolation and context window extension [32]. Length extrapolation techniques aimed to maintain stable PPL regardless of text length by designing specialized positional encodings or window attention mechanisms [6, 17, 18, 15, 14]. Conversely, context window extension methods focused on extending the context window of existing models by adapting positional encodings or temperature hyper-parameters, thereby enlarging the context window with minimal performance loss [13, 12, 16, 20, 11, 10]. This paper bridges the concepts of length extrapolation and context window extension through the lens of positional vectors, enhancing the interpretability of context windows in LLMs.
@@ 6 Conclusion
7 Limitation
Our work provides an extensive discussion and analysis of the context window through the lens of positional vectors. However, our study is mainly constrained by the use of small-scale LLMs that we trained ourselves, due to the unavailability of existing LLMs with the specific positional encodings and attention patterns required for our experiments. Though some mainstream LLMs are evaluated, these models are all based on RoPE. Furthermore, we have demonstrated the effectiveness of our proposed methods solely on our own models, again limited by the absence of suitable external models. In future work, we aim to seek a broader range of models to validate our findings more comprehensively.
Acknowledgement
This work was partially supported by National Natural Science Foundation of China under Grant No. 62222215, Beijing Municipal Science and Technology Project under Grant No. Z231100010323009, and Beijing Natural Science Foundation under Grant No. L233008. Xin Zhao is the corresponding author.
Appendix A Training and Experimental Details
@@
Appendix B The Positional Vectors After the First Layer
@@ B.1 Details
B.2 Proof of Preference in Attention Scores
@@
Appendix C Defination of Effective Interpolation Ratio
@@
Appendix D Analysis of Positional Vector Replacement
D.1 Optimal Replacement Layer
@@
D.2 Effective Interpolation Ratio
@@
Appendix E Pseudo Code
E.1 Positional Vector Replacement
@@
E.2 Attention Window Extension
@@
Appendix F Results of Additional LLMs
To ensure a fair comparison of positional vectors across various attention mechanisms and positional encodings, we conducted continual pre-training using TinyLlama under consistent settings. However, TinyLlama is a relatively small language model with suboptimal performance and continual pre-training may result in positional vectors exhibiting properties distinct from those obtained through training from scratch. Therefore, we selected three mainstream LLMs: Llama-3-8B [34], Yi-9B [35], and Qwen1.5-7B [36], for comparison. Additionally, we trained a new LLM, TL-NoPE-new, from scratch under the same conditions as TL-NoPE. In a similar vein, we extracted positional vectors using 32K samples from the RedPajama dataset.
F.1 Formation of Positional Vectors within Context Window
Figure 10: PCA visualization of positional vectors from the 1-st layer of Llama-3-8B, Qwen1.5-7B, Yi-9B, and TL-NoPE-new.
Through principal component analysis (PCA), we first visualize the positional vectors from the initial layer of these LLMs, as illustrated in Figure 10. Consistent with our expectations, the initial tokens exhibit distinct positional information, while the subsequent tokens display a high degree of similarity. This observation supports the conclusion that the first-layer attention mechanism makes the initial tokens form unique positional information, as discussed in Section 3.2.1.
Furthermore, we remove different components of the value vectors at different positions across all attention heads after the first layer. We then evaluate the impact of these modifications on both the positional vectors and the perplexity (PPL). As shown in Table 6, removing the positional basis of the initial tokens significantly degrades the model’s performance. Conversely, removing the components from subsequent tokens has a relatively smaller effect, highlighting the pivotal role of initial token positioning in influencing later tokens. However, we observe that larger LLMs are more attuned to semantic information and less affected by the removal of positional vectors compared to smaller LLMs.
Table 6: Results of removing different components in attention.
original w/o value w/o positional vector w/o positional basis w/o semantic basis
- 0$\sim$4 32-256 0$\sim$4 32$\sim$256 0$\sim$4 32-256 0$\sim$4 32-256
simi 1 0.75 0.9995 0.75 0.9583 0.2059 0.9997 0.9259 0.8666
Llama-3 ppl 6.74 16.27 6.63 17.20 8.4 >1000 6.60 17.6 15.18
simi 1 0.98 0.9999 0.92 0.9998 0.5368 1 0.91 0.9996
Yi-9B ppl 7.08 8.03 6.56 37.92 6.62 >1000 6.52 42.271 7.08
simi 1 0.98 0.9997 0.9847 0.9986 0.7382 0.9993 0.9998 0.9951
Qwen-1.5-7B ppl 7.97 9.51 8.03 9.51 8.04 217.13 7.98 8.09 8.68
simi 1 0.70 0.95 0.69 0.95 0.41 0.93 0.99 1.0
TL-NoPE-new ppl 11.03 224.74 22.36 263.53 20.91 >1000 21.78 11.66 12.699
F.2 Effect of Positional Vectors on Attention
In line with the experiments conducted in Section 3.2.3, we extract various components from the keys and queries to assess their impact on attention scores. The attention maps for the first 50 tokens are illustrated in Figure 11. When both the positional vectors and positional basis are removed, attention sinks disappear, and long-term decay is observed, which is consistent with the behavior seen in TinyLlama.
Figure 11: Logarithmic attention maps of Llama-3-8B, Qwen1.5-7B, Yi-9B, and new TL-NoPE.
F.3 Effect of Positional Vectors Beyond Context Window
Table 7: Resuls of PPL and change of positional vectors during direct extrapolation.
model context window(C) PPL(C) PPL(2C) Simi(2C)
Llama-3-8B 8192 6.74 >1000 0.30
Yi-9B 4096 7.08 102.58 0.24
TL-NoPE-new 2048 11.75 351.49 0.71
Table 8: Change of attention sinks and output logits beyond context window.
[…]
attention sink 0.467 0.1 0.005
Llama-3-8B 8192 logits similarity 1 0.9 0.88
attention sink 0.68 0.344 0.056
Yi-9B 4096 logits similarity 1 0.98 0.97
attention sink 0.17 0.02 0.0006
TL-NoPE-new 2048 logits similarity 1 0.97 0.9
We investigate the behavior of positional vectors under direct extrapolation scenarios. Specifically, we evaluate LLMs on sequences twice the length of their maximum context window, comparing the variations in PPL and positional vectors. The results, as shown in Table 7, indicate a sharp increase in PPL once the context window is exceeded. Additionally, the similarity between positional vectors within and beyond the context window decreases, highlighting the relationship between length extrapolation and the stability of positional representations.
Furthermore, we assess the impact of OOD positional vectors. We analyze the attention scores of initial tokens and the positional vectors of output logits within and beyond the context window. As demonstrated in Table 8, exceeding the context window results in the loss of the attention sinking property, and the positional vectors of the logits exhibit a different distribution compared to those within the context window. This suggests that OOD positional vectors influence the model’s distribution. Moreover, OOD positional encodings play a crucial role in large language models (LLMs) that leverage RoPE [15].
F.4 Experiments of Positional Vector Replacement
To assess whether our positional vector replacement method is specific to TL-NoPE, we applied it to TL-NoPE-new, which was trained from scratch. We evaluated its language modeling performance on the PG-19 dataset, using the same experimental setup as TL-NoPE. As shown in Table 9, our approach effectively extends the context window to 8K tokens and achieves performance comparable to attention scaling.
Table 9: Results of language modeling in PG-19 with TL-NoPE-new
Interpolation Method Factor 2048 4096 6144 8192
- - 29.0 110.7 634.8 1078.5
1.2 29.0 27.3 - -
Attention Scaling 1.5 36.8 41.6 44.4 46.2
r=2.5,$\alpha$=0.8 32.8 29.5 - -
Positional Vector Replacement(ours) r=5, $\alpha$=0.8 57.4 50.8 40.0 44.2